GRADUATE STUDENT COURSE ASSIGNMENTS

Block 0 Analysis —- Fefferman  Week 1
Block 1 Discrete Mathematics —- Babai  Weeks 1–8
Block 2a Problems —- Abert  Weeks 1–4
Block 2b Hyperplanes —- Klivans  Weeks 5–6
Block 2c A piacere —- May  Weeks 7–8
Block 3a Hyperbolic —- Muchnik and Bader  Weeks 1–4
Block 3b Mapping class groups —- Farb and Hruska  Weeks 5–8
Block 4 Applied mathematics —- Cattaneo Gordon Kirr Lewicka Weeks 1–8
Block 5 Apprentice Program —- Abert Babai May  Weeks 1*–4
Week 1 Week 2 Week 3 Week 4 Week 5 Week 6 Week 7 Week 8
Block 1 Block 1 Block 1 Block 1 Block 1 Block 1 Block 1 Block 1
Block 2a Block 2a Block 2a Block 2a Block 2b Block 2b Block 2c Block 2c
Block 3a Block 3a Block 3a Block 3a Block 3b Block 3b Block 3b Block 3b
Block 4 Block 4 Block 4 Block 4 Block 4 Block 4 Block 4 Block 4
Block 5* Block 5 Block 5 Block 5

*No class meetings first week, but meetings with students

Block 1 (Babai) (weeks 1–8) Chakraborty, Ganapathy, Raghunathan, Purdy, Schedler

Block 2a (Abert) (weeks 1–4) Balduzzi, Copeland, Cudney, Krishnan, Patel, Wallace

Block 2b (Klivans) (weeks 5–6) Balduzzi, Bremer, Copeland, Cudney, Patterson, Wallace

Block 2c (May) (weeks 7–8) Everyone is welcome to join me at lunch, and to talk with interested students afterwords. I’ll give an informal talk every lunch hour.

Block 3a (Muchnik and Bader) (weeks 1–4) Day, Dymarz, Johns, Ponto, Vasilyev

Block 3b (Farb and Hruska) (weeks 5–8) Bremer, Dymarz, Johns, Krishnan, Ponto, Vassilyev

Block 4 (Cattaneo Gordon Kirr Lewicka Lewicka) (weeks 1–8) Bremer (7–8), Carstea, Gurski, Morris, Rule

Block 5 (Apprentice) (weeks 1–4, possibly 5–8) Abouzaid, Johnson (3–8), Lee, Noel, Patterson, Ren

2 GRADUATE STUDENT COURSE ASSIGNMENTS

Assignments by block and by mentoring group

Name  Block  Weeks  Mentoring
Mohammed Abouzaid  5  4 weeks (first 4)  Apprentice
David Balduzzi  2  8 weeks  YSP
Chris Bremer  2b, 4d  4 weeks (last 4)  SESAME
Catalin Carstea  0 and 4  8 weeks  SESAME
Jeremy Copeland  2  8 weeks  YSP
Richard Cudney  2  8 weeks (miss 1st?)  YSP
Matthew Day  3a  4 weeks (first 4)  SESAME
Tullia Dymarz  3  6-7 weeks (not 6 and/or 7)  YSP
Nick Gurski  4  8 weeks  SESAME
Niles Johnson  2 and/or 5  6 weeks (3rd to 8th)  Apprentice
Joseph Johns  3  8 weeks  YSP
Sanjeevi Krishnan  2a, 3b  8 weeks  YSP
Ben Lee  5  4 weeks (first 4)  Apprentice
Courtney Morris  4  8 weeks  SESAME
Justin Noel  2  8 weeks  Apprentice
Deepam Patel  2a  4 weeks (first 4)  Apprentice
Eric Patterson  2b and 5  8 weeks  Apprentice
Kate Ponto  3  8 weeks  SESAME
Wei Ren  5  8 weeks  Apprentice
David Rule  0 and 4  8 weeks  SESAME
Travis Schedler  1  8 weeks  —-
Sergey Vasilyev  3  7 weeks (not 1st)  SESAME
Chris Wallace  2  8 weeks  YSP
Sourav Chakraborty  1  8 weeks  —-
Murali Ganapathy  1  8 weeks  —-
Nanda Raghunathan  1  8 weeks  —-
Eric Purdy (UG)  1  8 weeks  —-

ORGANIZERS

Team leaders for courses: people in charge of organizing problem sessions, making sure that graduate students are at talks, and coordinating with faculty, whose preferences on how graduate students are to help out are paramount.

1 Babai: Sourav Chakraborty

2a Abert: David Balduzzi and Chris Wallace 2b Klivans: Richard Cudney and Eric Patterson 3a Muchnik-Bader: Matthew Day and Tullia Dymarz 3b Farb-Hruska: Chris Bremer and Tullia Dymarz 4ab Cattaneo-Gordon: Catalin Carstea and David Rule 4cd Kirr-Lewicka: Nick Gurski and Courtney Morris

5 Abert-Babai-May: Justin Noel and Niles Johnson

Mentoring organizers: people in charge of setting up mentorship pairings.

APPRENTICE: Mohammed Abouzaid (Justin Noel and Niles Johnson weeks 5–8) Mentors: Abouzaid, Johnson*, Lee, Noel*, Patel, Patterson*, Ren* *Also YSP and SESAME as needed, especially during weeks 4–8.

YSP: David Copeland Mentors: Balduzzi, Copeland, Cudney, Dymarz, Krishnan, Wallace

SESAME: Kate Ponto Mentors: Bremer, Carstea, Day, Gurski, Morris, Ponto, Rule, Vasilyev